OpenStudy (anonymous):

Find an equation for the line tangent to the graph of f(x) = \frac{e^{6 x}}{8 x - 1} at the point (3, f(3)).

OpenStudy (anonymous):

\[Find an equation for the line tangent \to the graph of f(x) = \frac{e^{6 x}}{8 x - 1} at the point (3, f(3)). \]

OpenStudy (anonymous):

The slope at a point is determined by the derivative at that point. Find such derivative and use point-slope to create this line to be tangent to the graph.

OpenStudy (anonymous):

OKay so get the derivative and plug in the points for the slope?

OpenStudy (anonymous):

Yes, plug in \(x=3\) to the derivative for the slope, and then use point-slope formula: \[ y-y_0=m(x-x_0) \]To create such a line (where \(y_0=f(3), x_0=3\))

OpenStudy (anonymous):

Thanx

OpenStudy (anonymous):

The derivative is messy for m

OpenStudy (anonymous):

Hello!~~~~~~

OpenStudy (anonymous):

yes... I agree

OpenStudy (anonymous):

Did you get a nice complicated m once you plugged in your numbers?

OpenStudy (anonymous):

I got\[\frac{130e ^{18}}{529}\]

OpenStudy (anonymous):

but f(3) is equally complicated, so maybe the complications will cancel out in the equation for the line.

OpenStudy (anonymous):

yea and I dont know if anything is right

OpenStudy (anonymous):

so what did you get for the answer?

OpenStudy (anonymous):

f(3)=\[\frac{e ^{18}}{23}\]

OpenStudy (anonymous):

for the equation?

OpenStudy (anonymous):

For the line\[y- \frac{e ^{18}}{23}= \frac{130e ^{18}}{529}(x-3)\]

OpenStudy (anonymous):

in y=mx+b form

OpenStudy (anonymous):

Distribute, take the e term over to the right, find a common denominator and combine like terms.... what did you get? :)

OpenStudy (anonymous):

Stop making me learn just give me the answer. haha

OpenStudy (anonymous):

I'm mean that way. :)

OpenStudy (anonymous):

Did you read my testimonial for you?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

click on your icon, then click on it in the little window, then find testimonials on the left and click on the number

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

Your score was about 16 when we started today... look now... you have double that!

OpenStudy (anonymous):

yay!

OpenStudy (anonymous):

Now if only i could finish my homework

OpenStudy (anonymous):

k... now I gave you the point slope form for your problem... simplify it... you're in calculus... you can do it!

OpenStudy (anonymous):

Algebra is the hardest part of calculus.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

um........

OpenStudy (anonymous):

im trying to get b with this

OpenStudy (anonymous):

yeah, yeah, um.... I sat by your last post for 35 minutes before I realized that you had moved on to this question. I hope that you got the last one right. Yeah b is the hardest part.

OpenStudy (anonymous):

\[\frac{ e^{18} }{ 23}=\frac{ 130e ^{18} }{529}(3) +b\]

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

You did? Sorry, haha.

OpenStudy (anonymous):

My bad

OpenStudy (anonymous):

no biggie... this post is so long... my LaTeX stopped working... just a sec.

OpenStudy (anonymous):

nope... that's not it. haven't looked at @dumbcow 's link yet

OpenStudy (anonymous):

doh! that gives it away a little...

OpenStudy (anonymous):

alright... I'll do some more...

OpenStudy (anonymous):

ahhhhhhhhhhhhhhhhhhhhhhh

OpenStudy (dumbcow):

yeah sorry about that....anyway @monique.awesome , your equation is good just solve for b

OpenStudy (anonymous):

\[y= \frac{130e ^{18}}{529}x- \frac{390e ^{18}}{529}+ \frac\]\[y=\frac{130e ^{18}}{529}x - \frac{367e ^{18}}{529}\]I f your LaTeX is lagging like mine... I will at least force you to read LaTeX.... Learning!

OpenStudy (anonymous):

-\[-\frac{ 367 }{ 529}e ^{18}\]

OpenStudy (anonymous):

I was working on it!

OpenStudy (anonymous):

Hey look sI got b yay!

OpenStudy (anonymous):

woof!

OpenStudy (anonymous):

I think... was it -(367e^18)/529?

OpenStudy (anonymous):

Thats what I got now I have to figure out how to put that crap in my program haha

OpenStudy (anonymous):

good luck!

OpenStudy (anonymous):

Oh no! Now you will use Wolfram Alpha. Be careful! It's bad for your learning if abused.

OpenStudy (anonymous):

Wha?

OpenStudy (anonymous):

I figured it out!

OpenStudy (anonymous):

nothing....

OpenStudy (anonymous):

You truly deserve awesome in your name.

OpenStudy (anonymous):

Sorry I have to take ths part of calculus for my degree

OpenStudy (anonymous):

but my last calculus class was two semesters ago

OpenStudy (anonymous):

Aw thanx

OpenStudy (anonymous):

What!?! This isn't just for fun?

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

I wish

OpenStudy (anonymous):

Okay I post new question

OpenStudy (anonymous):

k